OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6

OpenStudy (anonymous):

first you want to isolate either x or y, which is easy to do in the first equation by subtraction 3y and getting x=-3y-1 then you can solve it by substitution

OpenStudy (texaschic101):

or multiply equation 1 by -2 and then add the equations...this should eliminate the x's and allow you to solve for y. Then sub what you got for y back into either equation and solve for x.

OpenStudy (anonymous):

You can use substitution method to solve this. Take any equation like x+3y=-1 for example. Isolate the variable x so (-1 - 3y) = x so since x= -1-3y plug that into the second equation. 2(-1-3y)+2y=6 -2-6y+2y=6 Simplify -4y-2=6 Now solve for y -4y=8 y=-2 Since y=-2 now you can solve x (x+3(-2))= -1 x+-6= -1 Subtract -6 on both sides and x= 5 So your answer is (5, -2)